OpenStudy (anonymous):

WILL FAN AND MEDAL DEEZ DAYZ In isosceles triangle ABC, BC (squiggly line) = CA and the measurement of B is 50. Find the measurement of A and C plz!

OpenStudy (wolf1728):

Angle B = 50° Angle A = Angle B Therefore, Angle C = 180 -50° -50° Angle C = 80°

OpenStudy (anonymous):

yewhaaaaa..??

OpenStudy (wolf1728):

Agree?

OpenStudy (anonymous):

plz draw and explain better dood I cant understand dis stuff...

OpenStudy (wolf1728):

It is an isosceles triangle. We are told that BC = CA What does squiggly line mean? This ~ ???

OpenStudy (anonymous):

yea....

OpenStudy (anonymous):

not to interrupt you or anything, but may u please speed up the process..

OpenStudy (anonymous):

...whaaa

OpenStudy (anonymous):

y it take u so long....

OpenStudy (anonymous):

#waytoolong

OpenStudy (wolf1728):

Okay BC = CA - it's an isoscleles triangle Angle B is 50 Therefore angle A = 50 it's an isoscleles triangle All triangles have 180° So we have 2 angles totalling 100° So the remaining angle is 80° Done

OpenStudy (wolf1728):

Sorry, couldn't find the actual '=~' symbol - wasted a LOT of time doing that because the puzzle didn't want to state things n plain English.

OpenStudy (anonymous):

ok, but I still don't actually understand. How did you get the angle of 50 degrees for A?

OpenStudy (wolf1728):

Angle A = Angle B = it is an isosceles triangle.

OpenStudy (wolf1728):

Thanks for the medal and fan - but you see how all the calculations are done right?

OpenStudy (anonymous):

ok I unserstand

OpenStudy (wolf1728):

okay - just making sure :-)